### v3

# Correlated Systematic Uncertainties and Errors-on-Errors in Measurement Combinations with an Application to the $7$-$8$ TeV ATLAS-CMS Top Quark Mass Combination

Abstract
The Gamma Variance Model (GVM) is a statistical model that incorporates uncertainties in the assignment of systematic errors (informally called errors-on-errors). The model is of particular use in analyses that combine the results of several measurements. In the past, combinations have been carried out using two alternative approaches: the Best Linear Unbiased Estimator (BLUE) method or what we will call the nuisance-parameter method. In this paper we derive useful relations that facilitate application of the model in practice. Specifically, we obtain a general relation between the BLUE and nuisance-parameter methods when the correlations induced by systematic uncertainties are non-trivial (i.e., not ±$\pm 1$ or 00), and we then generalise the nuisance-parameter approach to include errors-on-errors. We then present analytical formulas for estimating central values, confidence intervals, and goodness-of-fit when errors-on-errors are incorporated into the statistical model. To illustrate the properties of the GVM, we apply it to the 77778888 TeV ATLAS-CMS top quark mass combination. We also explore a hypothetical scenario by artificially adding a fictitious measurement as an outlier to the combination, illustrating a key feature of the GVM — its sensitivity to the internal consistency of the input data — which could become relevant for future combinations.
pacs:
02.50.TtInference methods and 02.70.RrGeneral statistical methods and 14.65.HaTop quarks
Keywords: Gamma Variance Model, errors-on-errors, combining measurements, correlated systematics, nuisance parameters, Bartlett correction, ATLAS-CMS top mass combination
1 Introduction
In Particle Physics analyses such as those at the Large Hadron Collider (LHC), many results are already dominated by systematic uncertainties or will be shortly. In this context, uncertainties in the assigned values of systematic errors themselves – informally called errors-on-errors – become increasingly important. Often, systematic uncertainties are estimated using real datasets of control measurements or Monte Carlo simulations. However, this is not always the case. Theoretical uncertainties often require ad hoc procedures for estimation, which can carry an intrinsic degree of uncertainty. Additionally, “two-point systematics”, where uncertainties are estimated from the difference in results from two methods, provide only a limited picture of the actual uncertainty. The Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 provides a statistical framework to handle errors-on-errors.
In this paper, we focus on the application of the GVM to combinations, and we derive a number of new results that facilitate use of the model in a non-trivial analysis. We first show how a likelihood with nuisance parameters can be constructed when the correlations induced by systematic uncertainties are not simply ±$\pm 1$ or 00, and how this approach relates to the Best Linear Unbiased Estimator (BLUE) method bib:Lyons ; bib:Valassi2003 ; bib:Nisius2020 ; bib:Nisius2014 . Next, we demonstrate how this model can be generalized to include errors-on-errors and derive analytical formulas for its application under these conditions. This framework is then applied to the $7$-$8$ TeV ATLAS-CMS top quark mass combination bib:ATLAS/CMS . Specifically, we investigate the impact on the combination’s central value and confidence interval when the largest sources of systematic uncertainties are treated as uncertain. We present the results by considering the largest systematic uncertainties as uncertain, one at a time, and we vary their associated error-on-error parameters. This approach serves as a general method to identify the systematic uncertainties to which a combination is sensitive when errors-on-errors are considered.
Finally, we explore how the inclusion of a hypothetical measurement that is in tension with the rest of the data affects the combination’s sensitivity to uncertain systematics. This exercise can be relevant for future combinations, as some existing top mass measurements deviate from the result of the ATLAS-CMS combined result bib:Tevatron2014 ; bib:ATLAS2023 . This example illustrates the behavior of the GVM in scenarios where the inputs of a combination are not entirely internally compatible. Unlike a standard combination, incorporating errors-on-errors makes the central value of the combination, here taken as the maximum likelihood estimator for the parameter of interest, more robust to discrepancies between input data values. Specifically, the central value becomes less sensitive to outliers. At the same time, the presence of discrepancies results in an increase in the confidence interval for the parameter of interest. This increase reflects the uncertainty stemming from the fact that the measurements in the dataset are not entirely compatible.
In Sec. 2, we present an overview of how a combination can be performed using nuisance parameters under non-trivial correlation assumptions, and how this method relates to the BLUE approach. Section 3 illustrates how to extend combinations to take errors-on-errors into account. In Sec. 4, we provide analytical formulas for profiling over nuisance parameters, confidence intervals, and goodness-of-fit when errors-on-errors are incorporated into the statistical model. In Sec. 5, we apply the GVM to the 7-8 TeV ATLAS-CMS top mass combination and explore the impact of introducing an outlier to the combination. A summary and conclusions are presented in Sec. 6.
2 Equivalence between nuisance-parameter and BLUE methods in combinations with non-trivial correlations
In Particle Physics, statistical data analysis aims at testing theoretical hypotheses $H$ using experimentally collected data, denoted as $\boldsymbol{y}$. The probability distribution of the data, $P(\boldsymbol{y}|H)$, is often indexed by a set of Parameters Of Interest (POIs) $\boldsymbol{\mu}$ and Nuisance Parameters (NPs) $\boldsymbol{\theta}$. NPs account for systematic uncertainties in the model, while the POIs are the main objective of the analysis. The likelihood function, $L(\boldsymbol{\mu},\boldsymbol{\theta})=P(\boldsymbol{y}|\boldsymbol{\mu},\boldsymbol{\theta})$, is the central object needed for inference about the POIs.
Here, we explore the application of combining $N$ measurements $\boldsymbol{y}=(y_{1},\ldots,y_{N})$ of a parameter, each with its own statistical and systematic uncertainties, to obtain a single best estimate. We assume that $\boldsymbol{y}$ follows a multivariate Gaussian distribution with expected values and covariance matrix given by
$\text{E}[y_{i}]$ $=$ $\mu+\sum_{s=1}^{M}\Gamma_{i}^{s}\theta_{s}\,,$ (1)
$V_{ij}$ $=$ $\text{cov}[y_{i},y_{j}]\,,$ (2)
where $M$ is the total number of NPs. Each NP $\theta_{s}$ represents a potential source of bias in the measured $y_{i}$ values, and the magnitude of these biases is described by the factors $\Gamma_{i}^{s}$. Following this notation, we will use indices $i,j$, etc., to refer to components of $\boldsymbol{y}$ and $s$ to label a source of systematic uncertainty.
Additionally, independent control measurements $\boldsymbol{u}$ are used to constrain the NPs. These control measurements $\boldsymbol{u}=(u_{1},\ldots,u_{M})$ are assumed to be best estimates of the NPs $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{M})$ and are treated as independently Gaussian-distributed random variables with standard deviations $\boldsymbol{\sigma_{u}}=(\sigma_{u_{1}},\ldots,\sigma_{u_{M}})$. Under these assumptions, the log-likelihood of the model is
$\begin{split}\ell(\mu,\boldsymbol{\theta})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta_{s}\right)-\frac{1}{2}\sum_{s=1}^{M}\frac{(u_{s}-\theta_{s})^{2}}{\sigma_{u_{s}}^{2}}\,.\end{split}$ (3)
It is usually convenient to redefine the factors $\Gamma_{i}^{s}$, the NPs $\theta_{s}$ and the auxiliary measurements $u_{s}$ such that $u_{s}$ is Gaussian distributed about $\theta_{s}$ with a variance of unity, i.e., $\sigma_{u_{s}}^{2}=1$. In the real experiment, the best estimates of the $u_{s}$ are zero. The idea here is that if there are known biases in the measurements, these are already subtracted. The remaining systematic uncertainty lies in the fact that this bias removal might be imprecise, due to the limited accuracy with which we can identify and quantify the biases. Nevertheless, here we will retain the complete notation in our discussion for greater generality. Auxiliary measurements must be explicitly included to perform Monte Carlo (MC) simulations, while the systematic variances $\sigma^{2}_{u_{s}}$ are needed to generalise the model to include errors-on-errors.
The above likelihood can be related to the Best Linear Unbiased Estimate (BLUE) method bib:Lyons ; bib:Valassi2003 ; bib:Nisius2020 ; bib:Nisius2014 , which is a commonly used technique for performing combinations that does not require the introduction of NPs. This method involves constructing an approximate Gaussian likelihood to describe the $N$ measurements of the combination, $\boldsymbol{y}$, based on a covariance matrix $W_{ij}$ that includes both statistical and systematic uncertainties. The resulting log-likelihood
$\ell(\mu)=-\frac{1}{2}\sum_{i,j=1}^{N}(y_{i}-\mu)\,W^{-1}_{ij}\,(y_{j}-\mu)$ (4)
is then maximized to get the best estimate of $\mu$. The fundamental assumption of the method is that the covariance matrix $W$ can be expressed as a sum of terms: one for the statistical uncertainties $V$ and one for each systematic source $U^{(s)}$:
$W_{ij}=V_{ij}+\sum_{s=1}^{M}U^{(s)}_{ij}\,.$ (5)
Here the index $s$ runs over each source of systematic uncertainty. These systematics are independent from each other in the sense that their associated control measurements are uncorrelated, but they induce correlations between the $\boldsymbol{y}$ measurements. This underlying structure is more evident when Eq.(3) is used instead. In fact, the two methods are equivalent when the auxiliary measurements are set to zero, $u_{s}=0$. Specifically, by profiling Eq.(3) over all the NPs, one can recover the BLUE log-likelihood as defined in Eq.(4). [^1] Profiling over a parameter means maximizing the likelihood with respect to that parameter while keeping all others fixed. Furthermore, the relationship between the factors $\Gamma_{i}^{s}$ and the BLUE covariance matrix $W$ uses Eq. (5) with
$U^{(s)}_{ij}=\Gamma_{i}^{s}\,\Gamma_{j}^{s}\,\sigma_{u_{s}}^{2}\,.$ (6)
The equivalence between these two methods is widely discussed in the literature; more details can be found, for example, in bib:Pinto2024 ; bib:Demortier1999 ; bib:Fogli2002 ; bib:Stump2001 ; bib:Thorne2002 ; bib:Botje2002 ; bib:Glazov2005 ; bib:Barlow2021 ; bib:List2010 ; bib:Aad2014 .
2.1 Systematics with non-trivial correlations
A limitation of the NP approach is that each systematic effect $s$ can only induce correlation coefficients of ±$\pm 1$ or 00, which we call ”trivial” correlations:
$\rho^{(s)}_{ij}=\frac{U^{(s)}_{ij}}{\sqrt{U^{(s)}_{ii}U^{(s)}_{jj}}}=\pm 1\text{ or }0\,.$ (7)
This restriction arises from the definition of the systematic terms $U^{(s)}_{ij}$ in the BLUE covariance matrix, as specified in Eq. (6). Generalising the BLUE method to accommodate this is very simple as it only requires modifying Eq. (6) to be
$U^{(s)}_{ij}=\rho^{(s)}_{ij}\Gamma_{i}^{s}\Gamma_{j}^{s}\sigma_{u_{s}}^{2}\,.$ (8)
Note that one may not, however, assign arbitrary values to the $\rho^{(s)}_{ij}$ as this could lead to $U^{(s)}$ not being positive definite.
The BLUE method has been used, e.g., in the top-mass combination bib:ATLAS/CMS that we consider in Sec. 5. To include errors-on-errors in it, we need to first relate the BLUE likelihood of the published analysis, which contains non-trivial correlations, to an equivalent formulation in terms of nuisance parameters. It is conceptually more complicated to extend the log-likelihood containing NPs of Eq. (3) so that it gives non-trivial correlations. To accomplish this, one needs to define a distinct NP $\theta^{i}_{s}$ associated with the systematic effect $s$ for each measurement $y_{i}$, and introduce correlations between the corresponding auxiliary measurements such that
$\mbox{cov}[u^{i}_{s},u^{j}_{s}]=\sigma_{u_{s}}^{2}\rho^{(s)}_{ij}\,.$ (9)
Thus the log-likelihood of Eq. (3) becomes
$\begin{split}\ell(\mu,\boldsymbol{\theta})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta^{i}_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta^{j}_{s}\right)\\ &-\frac{1}{2}\sum_{s=1}^{M}\sum_{i,j=1}^{N}(u^{i}_{s}-\theta^{i}_{s})\frac{1}{\sigma_{u_{s}}^{2}}\left(\rho^{(s)}\right)_{ij}^{-1}(u^{j}_{s}-\theta^{j}_{s})\,.\end{split}$ (10)
In App. A we prove the important result that this log-likelihood becomes equivalent to the BLUE method when the auxiliary measurements $u^{i}_{s}$ are set to zero. This equivalence can be demonstrated by profiling Eq.(10) with respect to all the NPs, which leads to the BLUE likelihood as defined in Eq.(4), with covariance matrix specified by Eqs. (5) and (8).
Sometimes, given a systematic effect $s$, one does not need to define an individual NP $\theta^{i}_{s}$ for each measurement $i$, as many induced correlations will still be ±$\pm 1$ or 00. For example, one may have two groups of measurements A and B, and a systematic effect s which is fully correlated inside each group of measurements, but induces a $0.5$ correlation in between the two groups. In this case one would only need to introduce two NPs for the systematic effect $s$, $\theta^{A}_{s}$ and $\theta^{B}_{s}$.
Note that here we are not introducing correlations between estimates of NPs associated with different systematics, as is done in the model proposed in bib:Kieseler2017 . That model addresses fits with simultaneously constrained uncertainties, where estimates of POIs and NPs are correlated with each other. Here, instead, for each systematic $s$, we introduce a separate NP $\theta^{i}_{s}$ and auxiliary measurement $u_{s}^{i}$ for each measurement $i$, and we correlate only the auxiliary measurements for a given systematic source $s$. We do this to match the BLUE result obtained with the systematic covariance matrix defined in Eq. (8).
3 Extending combinations to account for errors-on-errors
In this section we show how to extend a combined analysis of measurements to include errors-on-errors. We first treat in Sec. 3.1 the case where the systematic uncertainties relate to those of the corresponding BLUE analysis containing only “trivial” correlations, i.e., the correlation coefficients $\rho_{ij}^{(s)}$ in the systematic covariance matrix (8) are either 00 or ±$\pm 1$. Then in Sec. 3.2 we generalise the model to cover non-trivial correlations.
3.1 The GVM with trivial correlations
The likelihoods of Eqs. (3) and (10) can be extended to account for the presence of uncertainties in the assignment of systematic errors, informally denoted as errors-on-errors. This is achieved using the Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 . In this model the variances $\sigma^{2}_{u_{s}}$ of the control measurements $u_{s}$ are treated as adjustable parameters, and each $\sigma^{2}_{u_{s}}$ is constrained by its best estimate $v_{s}=s_{s}^{2}$. The estimates $\boldsymbol{v}$ are assumed to follow independent gamma distributions, which are suitable for positive-valued random variables. Detailed methodology and motivation for this approach can be found in bib:Cowan2019 . The resulting log-likelihood of this model is:
$\begin{split}\ell(\mu,\boldsymbol{\theta},\boldsymbol{\sigma_{u}^{2}})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta_{s}\right)-\frac{1}{2}\sum_{s=1}^{M}\frac{(u_{s}-\theta_{s})^{2}}{\sigma^{2}_{u_{s}}}\\ &-\frac{1}{2}\sum_{s=1}^{M}\bigg{[}\left(1+\frac{1}{2\varepsilon^{2}_{s}}\right)\log{\sigma^{2}_{u_{s}}}+\frac{v_{s}}{2\varepsilon_{s}^{2}\sigma^{2}_{u_{s}}}\bigg{]}\,.\end{split}$ (11)
This likelihood can be maximized in closed-form with respect to $\sigma^{2}_{u_{s}}$ for fixed $\mu$ and $\boldsymbol{\theta}$ using
$\widehat{\widehat{\sigma_{u_{s}}^{2}}}=\frac{v_{s}+2\varepsilon_{s}(u_{s}-\theta_{s})^{2}}{1+2\varepsilon_{s}^{2}}\,.$ (12)
This yields the profile log-likelihood $\ell_{p}(\mu,\boldsymbol{\theta})=\ell(\mu,\boldsymbol{\theta},\boldsymbol{\widehat{\widehat{\sigma_{u}^{2}}}})$
$\begin{split}\ell_{p}(\mu,\boldsymbol{\theta})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta_{s}\right)\\ &-\frac{1}{2}\sum_{s=1}^{M}\left(1+\frac{1}{2\varepsilon_{s}^{2}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u_{s}-\theta_{s})^{2}}{v_{s}}\right]\,.\end{split}$ (13)
The parameter $\varepsilon_{s}$ quantifies the relative uncertainty associated with the systematic error $\sigma_{u_{s}}$ and is commonly referred to as the error-on-error parameter. For instance, if $\varepsilon_{s}$ equals $0.3$, this indicates that $s_{s}=\sqrt{v_{s}}$ has a 30% relative uncertainty as an estimate of $\sigma_{u_{s}}$. In most cases $\varepsilon_{s}$ parameters must be supplied by the analyst using whatever information is available on the uncertainties of $\sigma_{u_{s}}$. In the absence of the required expert knowledge to inform such a choice, results of the GVM may be presented as a function of assumed values for $\varepsilon_{s}$. The model has also been discussed in bib:Reader2024 in the context of Parton Distribution Functions (PDF) fitting.
3.2 The GVM with non-trivial correlations
The likelihood of Eq.(10) is designed to admit non-trivial correlations, i.e., not only 00 or ±$\pm 1$. Generalising it to include errors-on-errors requires extending the GVM beyond what was originally presented in bib:Cowan2019 . The method to achieve this is nevertheless similar to what has already been discussed: the variance $\sigma_{u_{s}}^{2}$ in Eq. (10) is treated as an adjustable parameter of the model, and $v_{s}$, its best estimate, is considered a gamma-distributed random variable. The resulting generalised log-likelihood is
$\begin{split}\ell(\mu,\boldsymbol{\theta},\boldsymbol{\sigma_{u}^{2}})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta^{i}_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta^{j}_{s}\right)\\ &-\frac{1}{2}\sum_{s=1}^{M}\sum_{i,j=1}^{N}(u^{i}_{s}-\theta^{i}_{s})\frac{1}{\sigma_{u_{s}}^{2}}\left(\rho^{(s)}\right)_{ij}^{-1}(u^{j}_{s}-\theta^{j}_{s})\\ &-\frac{1}{2}\sum_{s=1}^{M}\left[\left(N+\frac{1}{2\varepsilon^{2}_{s}}\right)\log{\sigma^{2}_{u_{s}}}+\frac{v_{s}}{2\varepsilon_{s}^{2}\sigma^{2}_{u_{s}}}\right]\,.\end{split}$ (14)
This likelihood can be maximized with respect to $\sigma^{2}_{u_{s}}$ using
$\widehat{\widehat{\sigma_{u_{s}}^{2}}}=\frac{v_{s}+2\varepsilon_{s}^{2}\sum_{i,j}^{N}(u_{s}^{i}-\theta_{s}^{i})\left(\rho^{(s)}\right)_{ij}^{-1}(u_{s}^{j}-\theta_{s}^{j})}{1+2N\varepsilon_{s}^{2}}\,,$ (15)
to obtain the profile log-likelihood:
$\begin{split}\ell_{p}(\mu,\boldsymbol{\theta})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\mu-\sum_{s=1}^{M}\Gamma_{i}^{s}\theta^{i}_{s}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\mu-\sum_{s=1}^{M}\Gamma_{j}^{s}\theta^{j}_{s}\right)\\ &-\frac{1}{2}\sum_{s=1}^{M}\left(N+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+\frac{2\varepsilon_{s}^{2}}{v_{s}}\sum_{i,j=1}^{N}(u^{i}_{s}-\theta^{i}_{s})\right.\\ &\left.\times\left(\rho^{(s)}\right)_{ij}^{-1}(u^{j}_{s}-\theta^{j}_{s})\right]\,.\end{split}$ (16)
Often, in a manner similar to that of Sec. 2, it is useful to redefine $\Gamma_{s}$, $\theta_{s}$ and $u_{s}$ such that if there are no errors-on-errors ($\sigma_{u_{s}}^{2}$ exactly known), then $u_{s}$ is Gaussian distributed with a mean $\theta_{s}$ and unit variance. In this way the best estimates of the variances are all $v_{s}=1$. This does not alter the approach used to include errors-on-errors, as it is merely a redefinition of parameters.
In practical applications, the most feasible approach to account for errors-on-errors involves using NPs exclusively for systematic sources where $\varepsilon_{s}\neq 0$, and profiling overall NPs introduced as standard quadratic constraints. This leads to a mixed likelihood with a quadratic part constructed using the BLUE approach, and with NPs introduced only for systematic sources with uncertain errors. This can be achieved by replacing $V$ with a BLUE covariance matrix $W$, as defined in Eq. 5, but summing only over the non-uncertain systematic errors while keeping in the likelihood only those NPs associated with uncertain systematics. This can be done in a similar manner to the proof of the equivalence between the NPs approach and the BLUE approach, as shown in App. A.
Notice that when the off-diagonal correlation coefficients $\rho^{(s)}_{ij}$ of Eq. (9) are set to zero for some systematic effect $s$, the logarithmic term for that systematic effect reduces to
$\left(N+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+2\varepsilon_{s}^{2}\sum_{i=1}^{N}\frac{(u^{i}_{s}-\theta^{i}_{s})^{2}}{v_{s}}\right]\,.$ (17)
This result stems from the assumption that all $u_{s}^{i}$ for a fixed $s$ share the same variance $\sigma_{u_{s}}^{2}$, with a single estimated value $v_{s}$. In practice, this means assuming that if one systematic uncertainty is overestimated or underestimated by a certain factor, this factor will be the same for all the measurements affected by that systematic. Therefore, this model is suitable for modeling systematic uncertainties of the same type or those estimated using similar procedures.
If one wants to consider both the auxiliary measurements as uncorrelated and the estimates of the systematic uncertainties as independent, one would define the logarithmic terms as
$\sum_{i=1}^{N}\left(1+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u^{i}_{s}-\theta^{i}_{s})^{2}}{v^{i}_{s}}\right]\,.$ (18)
In this case, the original GVM, as defined in Eq. (13), is recovered, with the systematic uncertainty $s$ treated as independent for each measurement $i$.
4 Useful formulas for practical implementation of errors-on-errors
In this section we present some useful formulas to simplify the implementation of uncertain systematic errors in practice. Although the information here is somewhat technical, it can be used to overcome what would otherwise be important obstacles that stem from the logarithmic rather than quadratic constraints that appear in GVM likelihoods. In this section, we focus on the GVM with trivial correlations. The case of the GVM with non-trivial correlations is addressed in App. B.
4.1 Profiled values of nuisance parameters
In statistical modeling, NPs are introduced to account for systematic uncertainties. To reduce the dimensionality of the parameter space and facilitate the optimization of the likelihood, it is common to compute the profiled values of the NPs. These are the values of the NPs that maximize the likelihood for a fixed value of the POI, $\mu$. We denote such profiled values as $\hat{\hat{\theta}}_{s}$. Here, we focus on computing $\hat{\hat{\theta}}_{s}$ in the context of the GVM which includes logarithmic terms in the likelihood that make closed-form maximization challenging.
To compute the profiled values of the NPs, we maximize the log-likelihood defined in Eq. (13) with respect to the NPs, while treating the POI as fixed. This involves solving the score equations $\partial\ell_{p}/\partial\theta_{s}=0$, leading to the following system of equations:
$\begin{split}\sum_{i,j=1}^{N}&\Gamma_{i}^{s}V_{ij}^{-1}\left(y_{j}-\sum_{p=1}^{M}\Gamma_{j}^{p}\hat{\hat{\theta}}_{p}\right)\\ &+\frac{1+2\varepsilon_{s}^{2}}{v_{s}+2\varepsilon_{s}^{2}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)^{2}}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)=0\,.\end{split}$ (19)
The score equations for the NPs form a system of coupled cubic equations, which do not have a general analytical solution. While it is always possible to perform a numerical optimization, it is also possible to solve them perturbatively at all orders in $\varepsilon_{s}^{2}$, using a recursive approach. This can be especially useful for complicated fits with many NPs.
These nonlinear equations can be solved using a perturbative expansion in powers of $\varepsilon_{s}^{2}$. We express $\hat{\hat{\theta}}_{s}$ as:
$\hat{\hat{\theta}}_{s}=\hat{\hat{\theta}}_{s}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{s}^{(1)}+\varepsilon_{s}^{4}\hat{\hat{\theta}}_{s}^{(2)}+\cdots\,,$ (20)
where $\hat{\hat{\theta}}_{s}^{(0)}$ is the solution when $\varepsilon_{s}^{2}=0$, corresponding to the case without errors-on-errors. Substituting this expansion into Eq. (19) and equating terms of equal powers of $\varepsilon_{s}^{2}$, we obtain a recursive set of equations for the coefficients $\hat{\hat{\theta}}_{s}^{(n)}$. Using this expansion, the first term in the $\hat{\hat{\theta}}_{s}$ series is given by
$\hat{\hat{\theta}}_{s}^{(0)}=\sum_{p=1}^{M}\left(C^{(0)}\right)_{sp}^{-1}\left[\sum_{i,j=1}^{N}\Gamma_{i}^{p}V_{ij}^{-1}\left(y_{j}-\mu\right)+\frac{u_{p}}{v_{p}}\right]\,,$ (21)
where the matrix $C^{(0)}_{sp}$ is defined as
$C^{(0)}_{sp}=\sum_{i,j=1}^{N}\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{v_{s}}\,,$ (22)
where $\delta_{sp}$ is the Kronecker delta. Notice that this result is also the solution to the quadratic likelihood of Eq. (3). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative correction factor $\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{s}^{(n)}$ will be
$\begin{split}\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{s}^{(n)}&=\sum_{p=1}^{M}\left(C^{(n)}\right)_{sp}^{-1}\left[\sum_{i,j=1}^{N}\Gamma_{i}^{p}V_{ij}^{-1}\bigg{(}y_{j}-\mu\right.\\ &\left.-\sum_{p^{\prime}=1}^{M}\Gamma_{j}^{p^{\prime}}T_{p^{\prime},n-1}\bigg{)}+\frac{u_{p}-T_{p,n-1}}{S_{p,n-1}^{2}}\right]\,,\end{split}$ (23)
where the matrix $C^{(n)}_{sp}$ is defined as
$C^{(n)}_{sp}=\sum_{i,j=1}^{N}\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{S_{s,n-1}^{2}}\,,$ (24)
while $T_{s,n}$ and $S_{s,n}^{2}$ are given by
$T_{s,n}=\hat{\hat{\theta}}_{s}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{s}^{(1)}+\cdots+\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{s}^{(n)}\,,$ (25)
$S_{s,n}^{2}=\frac{v_{s}+2\varepsilon_{s}^{2}\left(u_{s}-T_{s,n}\right)^{2}}{1+2\varepsilon_{s}^{2}}\,.$ (26)
In summary, $T_{s,n}$ and $S_{s,n}^{2}$ represent the order $\varepsilon_{s}^{2n}$ approximations of $\hat{\hat{\theta}}_{s}$ and $\widehat{\widehat{\sigma_{u_{s}}^{2}}}$, respectively.
The convergence of the perturbative expansions derived in Eqs. (20) requires the condition
$\frac{2\varepsilon_{s}^{2}}{v_{s}}(u_{s}-\hat{\hat{\theta}}_{s})^{2}<1$ (27)
to hold, defining the radius of convergence for the series. While the expansion often approximates the exact solution well even if this condition is slightly violated, it typically breaks down for large $\varepsilon_{s}^{2}$ when the cubic score equation admits three real solutions instead of one.
Using the perturbative solutions for $\hat{\hat{\theta}}_{s}$, the log-likelihood can be optimized with respect to the POI $\mu$ to determine its best estimate.
4.2 Confidence intervals
In frequentist statistics, a confidence region in the full parameter space $(\mu,\boldsymbol{\theta})$ can be constructed by testing hypothesized parameter values using the likelihood-ratio test statistic,
$w_{\mu\boldsymbol{\theta}}=-2\log\left(\frac{L(\mu,\boldsymbol{\theta})}{L(\hat{\mu},\hat{\boldsymbol{\theta}})}\right)=2\left[\ell(\hat{\mu},\hat{\boldsymbol{\theta}})-\ell(\mu,\boldsymbol{\theta})\right]\,.$
When constructing confidence intervals for the POI $\mu$ only, we use the profile likelihood ratio:
$w_{\mu}=-2\log\left(\frac{L(\mu,\hat{\hat{\boldsymbol{\theta}}})}{L(\hat{\mu},\hat{\boldsymbol{\theta}})}\right)=2\left[\ell(\hat{\mu},\hat{\boldsymbol{\theta}})-\ell(\mu,\hat{\hat{\boldsymbol{\theta}}})\right].$
Here, $\hat{\mu}$ and $\hat{\boldsymbol{\theta}}$ are the Maximum Likelihood Estimators (MLEs), which are the values of the POI and the NPs that maximize the likelihood, while $\hat{\hat{\boldsymbol{\theta}}}$ are the profiled values of the NPs, obtained by maximizing the likelihood with $\mu$ fixed.
Confidence intervals are typically computed under the assumption that the likelihood ratio $w_{\mu\boldsymbol{\theta}}$ asymptotically follows a $\chi^{2}$ distribution with $1+M$ degrees of freedom, where $M$ is the number of NPs in the model. This holds in the large sample (also called asymptotic) limit (see, e.g., bib:Cowan2011 ; bib:Algeri2020 ). In this limit, the profile likelihood ratio $w_{\mu}$ is $\chi^{2}$-distributed with 1 degree of freedom, corresponding to the single POI. This results in two equivalent methods for estimating confidence intervals: either by imposing $w_{\mu}<1$ or by requiring $\ell(\hat{\mu},\hat{\boldsymbol{\theta}})-\ell(\mu,\hat{\hat{\boldsymbol{\theta}}})<\frac{1}{2}$.
For Gaussian data with linear dependence on model parameters, these statistics are inherently $\chi^{2}$-distributed, making the asymptotic limit exact for Gaussian models defined by Eqs. (3) and (10). However, the GVM
log-likelihoods in Eqs. (11) and (14) contain non-quadratic terms, causing deviations from the $\chi^{2}$ distribution by terms of order $\varepsilon_{s}^{2}$, which affect the precision of confidence intervals. This issue can be addressed using the Bartlett correction bib:Bartlett1937 ; bib:Codeiro2014 . The Bartlett correction belongs to a part of statistics called higher-order asymptotics bib:Brazzale2007 , which addresses deviations from asymptotic limits in statistical models. A discussion on how to apply higher-order asymptotics to the GVM can be found in bib:Canonero2023 . Instead of redefining the likelihood ratio’s distribution, the Bartlett correction rescales it to better match a $\chi^{2}$ distribution, while ensuring $p$-values remain accurate up to terms of order $\varepsilon_{s}^{4}$.
Specifically, the Bartlett-corrected likelihood ratio is defined as
$w_{\mu\boldsymbol{\theta}}^{\ast}=w_{\mu\boldsymbol{\theta}}\,\left(\frac{M+1}{\text{E}[w_{\mu\boldsymbol{\theta}}]}\right)\equiv\frac{w_{\mu\boldsymbol{\theta}}}{1+b_{\mu\boldsymbol{\theta}}/(M+1)}\,,$ (28)
where $b_{\mu\boldsymbol{\theta}}=\text{E}[w_{\mu\boldsymbol{\theta}}]-1-M$ represents the correction to the asymptotic expectation value (which is $1+M$ for $w_{\mu\boldsymbol{\theta}}$). Equivalently, the Bartlett-corrected profile likelihood ratio is
$w_{\mu}^{\ast}=\frac{w_{\mu}}{\text{E}[w_{\mu}]}\equiv\frac{w_{\mu}}{1+b_{\mu}}\,,$ (29)
where $b_{\mu}=\text{E}[w_{\mu}]-1$. The Bartlett-corrected $w_{\mu\boldsymbol{\theta}}^{\ast}$ and $w_{\mu}^{\ast}$ follow $\chi^{2}$ distributions with reduced errors of order $\varepsilon_{s}^{4}$ (see bib:Canonero2023 ).
In the GVM likelihoods considered here, $\mu$ and $\boldsymbol{\theta}$ are location parameters, meaning that $b_{\mu\boldsymbol{\theta}}$ and $b_{\mu}$ are constants because the expectation values of $w_{\mu\boldsymbol{\theta}}$ and $w_{\mu}$ do not depend on $\mu$ or $\boldsymbol{\theta}$. Consequently, the methods described earlier are modified as follows.
The correction factors $b_{\mu\boldsymbol{\theta}}$ and $b_{\mu}$ can be estimated analytically to order $\varepsilon_{s}^{2}$. This perturbative calculation is carried out using the Lawley formula bib:Lawley1956 ; bib:Codeiro1993 . For the likelihood defined in Eq. (13), one can show that
$b_{\mu}=b_{\mu\boldsymbol{\theta}}-\tilde{b}_{\boldsymbol{\theta}}\,,$ (30)
following the methodology detailed in bib:Canonero2023 . The factors $b_{\mu\boldsymbol{\theta}}$ and $\tilde{b}_{\boldsymbol{\theta}}$ are found to be
$\begin{split}b_{\mu\boldsymbol{\theta}}&=\sum_{s=1}^{M}\left[4\frac{j^{\theta_{s}\theta_{s}}}{\widehat{\sigma_{u_{s}}^{2}}}-\left(\frac{j^{\theta_{s}\theta_{s}}}{\widehat{\sigma_{u_{s}}^{2}}}\right)^{2}\right]\varepsilon_{s}^{2}\,,\\ \tilde{b}_{\boldsymbol{\theta}}&=\sum_{s=1}^{M}\left[4\frac{\tilde{j}^{\theta_{s}\theta_{s}}}{\widehat{\sigma_{u_{s}}^{2}}}-\left(\frac{\tilde{j}^{\theta_{s}\theta_{s}}}{\widehat{\sigma_{u_{s}}^{2}}}\right)^{2}\right]\varepsilon_{s}^{2}\,.\end{split}$ (31)
Here, $\widehat{\sigma_{u_{s}}^{2}}$ denotes the MLE of $\sigma_{u_{s}}^{2}$, obtained by evaluating Eq. (12) at $\hat{\theta}_{s}$. The matrix $j^{-1}$ represents the covariance matrix of the MLEs derived from the original quadratic likelihood in Eq. (3). It is computed as:
$j^{-1}=\begin{pmatrix}j^{\mu\mu}&j^{\mu\boldsymbol{\theta}}\\[5.69046pt] j^{\boldsymbol{\theta}\mu}&j^{\boldsymbol{\theta}\boldsymbol{\theta}}\end{pmatrix}=-\left(\nabla^{2}_{\mu,\,\boldsymbol{\theta}}\,\ell\bigg{|}_{\hat{\mu},\,\hat{\boldsymbol{\theta}}}\right)^{-1},$ (32)
where $\nabla^{2}_{\mu,\,\boldsymbol{\theta}}\,\ell$ denotes the matrix of second derivatives of the log-likelihood with respect to all parameters $(\mu,\boldsymbol{\theta})$. Upper indices are used to label the components of $j^{-1}$. The matrix $\tilde{j}^{-1}$ is defined by treating $\mu$ as fixed,
$\tilde{j}=\begin{pmatrix}\tilde{j}^{\mu\mu}&\tilde{j}^{\mu\boldsymbol{\theta}}\\[5.69046pt] \tilde{j}^{\boldsymbol{\theta}\mu}&\tilde{j}^{\boldsymbol{\theta}\boldsymbol{\theta}}\end{pmatrix}=\left(-\nabla^{2}_{\boldsymbol{\theta}}\,\ell\bigg{|}_{\,\hat{\boldsymbol{\theta}}}\right)^{-1}\,,$ (33)
and it is also computed using the quadratic likelihood given in Eq. (3). More details on applying the Lawley formula to the GVM can be found in bib:Canonero2023 , where the Bartlett correction factor $b_{\mu}$ was found to be zero, as the parameters $\mu$ and $\boldsymbol{\theta}$ were treated as orthogonal, meaning that their estimators were considered uncorrelated, i.e., $j^{\theta_{s}\theta_{s}}=\tilde{j}^{\theta_{s}\theta_{s}}$. Equation (30) supersedes this result by providing a more precise correction.
These $\varepsilon_{s}^{2}$ approximations are highly effective when the condition specified in Eq. (27) is satisfied.
4.3 Goodness-of-fit
To evaluate how well the selected model fits the observed data, a Goodness-of-Fit (GOF) statistic can be used. This can be defined as
$q=-2\log\frac{L(\hat{\mu},\hat{\boldsymbol{\theta}})}{L_{\rm s}(\hat{\boldsymbol{\varphi}},\hat{\boldsymbol{\theta}})}\,,$ (34)
where $L_{\rm s}$ represents the likelihood of the saturated model. In the saturated model, the expected values $\text{E}[y_{i}]=\mu$ are replaced by independent parameters $\boldsymbol{\varphi}=(\varphi_{1},\dots,\varphi_{N})$, such that $\text{E}[y_{i}]=\varphi_{i}$. Since there is one free parameter $\varphi_{i}$ for each observed value $y_{i}$, the log-likelihood of the saturated model is zero for Eq. (13). Therefore, the GOF statistic simplifies to
$\begin{split}q=&\sum_{i,j=1}^{N}\left(y_{i}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{i}^{s}\hat{\theta}_{s}\right)V_{ij}^{-1}\left(y_{j}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{j}^{s}\hat{\theta}_{s}\right)\\ &+\sum_{s=1}^{M}\left(1+\frac{1}{2\varepsilon_{s}^{2}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u_{s}-\hat{\theta}_{s})^{2}}{v_{s}}\right]\end{split}$ (35)
for the GVM with trivial correlations.
To assess how well the model fits the observed data, one can compute a $p$-value from the GOF statistic. In the limit as $\varepsilon_{s}\to 0$, the GOF statistic reduces to a sum of squares and follows a $\chi^{2}$ pdf with $N-1$ degrees of freedom. The $p$-value can then be calculated as $p=1-F_{\chi^{2}}(q_{\rm obs})$, where $F_{\chi^{2}}$ is the cumulative distribution function of a $\chi^{2}$ distribution, and $q_{\rm obs}$ is the value of $q$ computed with the observed data. Another common metric to assess the quality of the fit is the ratio $q/N_{\text{dof}}$. If $q$ follows a chi-squared distribution with $N_{\text{dof}}$ degrees of freedom, its expected value is $N_{\text{dof}}$, with a standard deviation of $\sqrt{2N_{\text{dof}}}$. Therefore, if the model reasonably describes the data, this ratio is expected to be close to $1$.
For $\varepsilon_{s}>0$, the $\chi^{2}$ distribution approximates the true distribution, with an error of order $\varepsilon_{s}^{2}$, similar to the case of the likelihood ratio. To improve this approximation, we can apply a Bartlett correction to the GOF statistic, defined as
$q^{\ast}=q\,\frac{N-1}{E[q]}=\frac{q}{1+b_{q}/(N-1)},$ (36)
where $E[q]$ is the expected value of $q$ and $b_{q}=E[q]-N+1$. In the limit as $\varepsilon_{s}\to 0$, $b_{q}\to 0$, and thus $q^{\ast}$ converges to $q$. For values of $\varepsilon_{s}>0$, the corrected statistic $q^{\ast}$ follows a $\chi^{2}$ distribution with $N-1$ degrees of freedom, up to error terms of order $\varepsilon_{s}^{4}$. Therefore, the corrected $p$-value $p^{\ast}=1-F_{\chi^{2}}(q^{\ast}_{\rm obs})$ provides a more reliable measure of how well the model fits the data. Alternatively, one can use the corrected ratio $q^{\ast}/N_{\text{dof}}$ to assess the quality of the fit.
The factor $b_{q}$ can be computed analytically to order $\varepsilon_{s}^{2}$ by expressing the expectation value of $q$ as $\text{E}[q]=-2\text{E}[\ell(\mu,\boldsymbol{\theta})]-\text{E}[w_{\mu\boldsymbol{\theta}}]$. The term $2\text{E}[\ell(\mu,\boldsymbol{\theta})]$ is computed using equation $(3)$ from bib:Canonero2023 . For the model of Eq. (13), the result is
$b_{q}=3\sum_{s=1}^{M}\varepsilon_{s}^{2}-b_{\mu\boldsymbol{\theta}}\,,$ (37)
where $b_{\mu\boldsymbol{\theta}}$ is defined in Eq. (31).
5 Application to $7$-$8$ TeV ATLAS-CMS top quark mass combination
[…]
All-hadronic bib:ATLAS2014 7 $175.06\pm 1.35\pm 1.21$$\pm 1.82$
Dileptonic bib:ATLAS2015 7 $173.79\pm 0.54\pm 1.31$$\pm 1.42$
Lepton+jets bib:ATLAS2015 7 $172.33\pm 0.75\pm 1.04$$\pm 1.28$
All-hadronicbib:ATLAS2017 8 $173.72\pm 0.55\pm 1.02$$\pm 1.16$
Dileptonic bib:ATLAS2016 8 $172.99\pm 0.41\pm 0.74$$\pm 0.84$
Lepton+jets bib:ATLAS2018fwq 8 $172.08\pm 0.39\pm 0.82$$\pm 0.91$
Table 1: Top quark mass measurements in $t\bar{t}$ final states at different center-of-mass energies, based on ATLAS data
[…]
All-hadronic bib:CMS2014 7 $173.49\pm 0.69\pm 1.23$$\pm 1.41$
Dileptonic bib:CMS2012a 7 $172.50\pm 0.43\pm 1.52$$\pm 1.58$
Lepton+jets bib:CMS2012b 7 $173.49\pm 0.43\pm 0.97$$\pm 1.06$
All-hadronicbib:CMS2016a 8 $172.32\pm 0.25\pm 0.57$$\pm 0.62$
Dileptonic bib:CMS2016a 8 $172.22\pm 0.18\pm 0.94$$\pm 0.95$
Lepton+jets bib:CMS2016a 8 $172.35\pm 0.16\pm 0.45$$\pm 0.48$
Single top bib:CMS2017 8 $172.95\pm 0.77\pm 0.93$$\pm 1.20$
$J/\psi$ bib:CMS2016b 8 $173.50\pm 3.00\pm 0.94$$\pm 3.14$
Secondary vertex bib:CMS2016c 8 $173.68\pm 0.20\pm 1.11$$\pm 1.12$
Table 2: Top quark mass measurements in $t\bar{t}$ final states at different center-of-mass energies, based on CMS data
Figure 1: The plot shows the variation of the central value as a function of the error-on-error parameter $\varepsilon_{s}$. Each line represents the change of the central value when the systematic uncertainties in the legend are considered uncertain one at a time. The central values are computed explicitly at points marked by dots and linearly interpolated in between
Figure 2: The plot shows the variation of the $68.3\%$ confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$. Each line represents the change of the central value when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
Measurements and combinations of the top quark mass are a relevant testing ground for the errors-on-errors framework for several reasons. Firstly, these measurements are becoming increasingly dominated by systematic uncertainties, making a precise evaluation of them crucial for the correct interpretation of results. Furthermore, measurements of the top quark mass are often significantly impacted by QCD modeling uncertainties, especially those exploiting jets in the final state, which themselves can be quite uncertain. In addition, the GVM’s results are sensitive to the internal compatibility of the data, and some of the top mass measurements exhibit tensions between each-other; for example, the Tevatron bib:Tevatron2014 combination is slightly incompatible with the result of ATLAS-CMS one bib:ATLAS/CMS , as well as the run $2$ ATLAS top mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 .
We begin by replicating the results of the 7-8 TeV ATLAS-CMS top mass combination bib:ATLAS/CMS using the BLUE method, as described in Sec. 2. To construct the BLUE matrix, we use Eqs. (5) and (8), incorporating data from tables A.2 and A.3 in the appendices of bib:ATLAS/CMS (Arxiv version) and appendix B of bib:ATLAS/CMSb . Specifically, the statistical covariance matrix, which is diagonal, incorporates the reported statistical uncertainties for each measurement. Additionally, each term $\Gamma^{s}_{i}$ corresponds to the error on the $i$-th measurement attributed to systematic source $s$. As discussed previously, we set $\sigma^{2}_{u_{s}}=1$. The correlation coefficients $\rho^{(s)}_{ij}$ are determined using the correlation matrices from appendix B of bib:ATLAS/CMSb . The result found using the BLUE approach is
$m_{t}=172.51\pm 0.33\,\text{GeV}\,.$ (38)
Our result has a minor discrepancy of $0.01$ GeV from the central value reported in bib:ATLAS/CMS , but the confidence interval agrees with that reported in the paper. Discrepancies of this magnitude can be expected since all public results are rounded to the second decimal place.
5.1 GVM analysis for combined estimate of $m_{t}$
To demonstrate the consistency of the proposed framework, we first show that the model proposed to handle systematic correlations (Eq. (10)) replicates the BLUE results. This is achieved by taking the limit $\varepsilon_{s}\rightarrow 0$ in Eq. (16), as in this limit, Eqs. (10) and (16) are equivalent. At the same time, we study how adding errors-on-errors to some systematics affects the result of the combination. We focus on each main systematic uncertainty separately, assuming that only one systematic includes an associated error-on-error at a time. This approach allows one to isolate the impact of one specific uncertain systematic error on the combination. We then vary the value of the error-on-error parameter for that systematic to determine the outcomes different values would produce.
To construct the profile likelihood for the GVM, we use Eq. (16). Furthermore, we maximize the log-likelihood with respect to the NPs using the perturbative approach described in Sec. 4.1, while employing the formulas provided in App. B to account for non-trivial correlations. NPs are introduced only for the systematic sources treated as uncertain, while all other systematic effects are handled using the BLUE approach (see Sec. 2). We set $u^{i}_{s}=0$ and $v_{s}=1$, so that the estimated systematic uncertainties are absorbed into the matrix $\Gamma$. To construct $\Gamma$ we use the uncertainties from tables A.2 and A.3 from the Arxiv version of bib:ATLAS/CMS , while the covariance matrices $\rho^{(s)}$ are derived from the correlation matrices provided in appendix B of bib:ATLAS/CMSb . Some of these correlation matrices exhibit one or more negative eigenvalues. We address this issue by adding a constant term equal to the smallest negative eigenvalue to the diagonal of each covariance matrix. This adjustment ensures that all correlation matrices are positive semi-definite. While this has a negligible effect on the overall combination, it is a critical step to ensure the convergence of the log-likelihood maximization process. This regularization step was not required for the BLUE approach, as the full covariance matrix $W$ in the BLUE method is positive semi-definite.
The ATLAS inputs to the combination bib:ATLAS2018fwq ; bib:ATLAS2015 ; bib:ATLAS2016 ; bib:ATLAS2014 ; bib:ATLAS2017 are summarized in Tab. 1, while the CMS ones bib:CMS2014 ; bib:CMS2012a ; bib:CMS2012b ; bib:CMS2016a ; bib:CMS2017 ; bib:CMS2016b ; bib:CMS2016c are listed in Tab. 2. We consider as potentially uncertain the eight primary sources of systematic uncertainty as listed in table 2 of bib:ATLAS/CMS , as we expect that the largest systematic errors would have a more pronounced effect on the final result if they are themselves uncertain. This selection does not imply that we consider these systematics as necessarily uncertain; rather, we intend to demonstrate the potential impact of various assumptions on the results of the combination.
Figure 1 illustrates the variation in the central value (i.e., Maximum Likelihood Estimator) of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. Specifically, the plot shows the dependence of the central value $\hat{\mu}$ on the error-on-error parameter $\varepsilon_{s}$ for each systematic uncertainty. The nomenclature for the systematic components utilized here is the same used in bib:ATLAS/CMS , where a detailed explanation of what they describe can be found. The most important conclusion here is that the central value of the combination is quite robust to the presence of uncertain systematic errors. Specifically, the change in the central value remains always within $0.1$ GeV for the explored range of $\varepsilon_{s}$ values, which is well within the confidence interval of approximately $0.3$ GeV. Ultimately, all central values converge to the BLUE result as $\varepsilon_{s}$ approaches zero, meaning that the way we implement correlations in our model, as described in Sec. 2, is consistent. The reason why the b tagging and QCD radiation lines do not converge to the same point is due to the regularization of their associated correlation matrices, as previously described.
Similarly, Fig. 2 displays the half-width of the $68.3\%$ confidence interval as a function of $\varepsilon_{s}$. As for the central value, each line corresponds to the case where only one systematic uncertainty at a time is considered as itself uncertain. The confidence intervals were computed using the Bartlett corrections, as described in Sec. 4.2, while employing the formulas provided in App. B to account for non-trivial correlations.
Figure 2 confirms our previous conclusion: the combination is generally robust to the presence of uncertain systematic errors. However, an increase of about $10\%$ in the confidence interval is observed when the b-JES systematic uncertainty is known with a relative error of approximately $50\%$. The combination’s heightened sensitivity to uncertainties in this systematic error stems from the fact that it constitutes the largest uncertainty of the combination. Notice that the b-JES systematic uncertainty, modeling the flavor response of b-JES, is a two-point systematic, i.e., based on the difference obtained with two methods, and thus falls into the category of systematic sources that could plausibly be uncertain. The confidence interval remains stable if any other of the systematic uncertainties considered are taken as themselves uncertain. A negligible bias in the confidence interval size of the b tagging and QCD radiation uncertainties is present due to the regularization of the correspondent correlation matrices.
5.2 Sensitivity to outliers
Figure 3: The plot shows the variation of the central value as a function of the error-on-error parameter $\varepsilon_{s}$ when one fictitious measurement is included in the combination. Each line represents the change of the central value when the systematic uncertainties in the legend are considered uncertain one at a time. The central values are computed explicitly at points marked by dots and linearly interpolated in between
Figure 4: The plot shows the variation of the $68.3\%$ confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$ when one fictitious measurement is included in the combination. Each line represents the change in the confidence interval when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
The conclusions of the last section would be significantly impacted if any of the combination inputs exhibited tension with the rest of the measurements. It is not uncommon to have input values that exhibit a significant tension and it could possibly happen for an updated LHC-Tevatron combination, or for a LHC Run $2$ combination that includes the ATLAS top mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 . However, since extending the combination to include additional measurements would require knowledge of the correlations between them and those listed in Tabs. 1 and 2, we introduce a fictitious measurement into the dataset to illustrate the properties of the GVM in such a scenario. Specifically, we consider adding a measurement with a central value of $m_{t}^{\mathrm{NEW}}=174.5$ GeV, a statistical uncertainty of $0.4$ GeV, and a global systematic uncertainty of $0.5$ GeV. We assume that the systematic uncertainty of the new measurement is uncorrelated with the uncertainties of the other measurements.
If the fictitious measurement is added to the combination, without considering errors-on-errors, the result obtained using the BLUE approach is $m_{t}^{(1)}=172.91\pm 0.29\,\text{GeV}$. As anticipated, the inclusion of the new measurement shifts the central value of the combination towards higher values and results in a reduction of the confidence interval.
Figure 3 illustrates the variations in the central value of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. On top of the systematics considered in the last section, the NEW systematic uncertainty on the fictitious measurement is also considered. In contrast to observations made in the previous section, the inclusion of the outlier in the dataset makes the effect of errors-on-errors non-negligible. The central value significantly shifts when the NEW systematic uncertainty is considered uncertain, moving back to that of the original combination. This is an important property of the GVM: while typically the presence of an outlier would significantly pull the result of a combination, within the GVM framework as the error-on-error parameter increases, the outlier is assigned lesser weight within the combination. Consequently, the central value shifts back towards that of the original combination, thereby diminishing the outlier’s effect.
Similarly, Fig. 4 displays the variation in the half-size of the $68.3\%$ confidence interval as a function of $\varepsilon_{s}$. The plot shows how the inclusion of the fictitious measurement affects the combination. It demonstrates that when the NEW systematic uncertainty includes an associated error-on-error, the confidence interval can increase significantly, by up to 25%. This is the second relevant property of the GVM: The size of the confidence interval is sensitive to the internal consistency of the input data. Specifically, the less compatible the input measurements are, the more the confidence interval is subject to inflation when errors-on-errors are considered. This is because the GVM treats the tension in the dataset as an additional source of uncertainty resulting in an inflated confidence interval.
The perturbative analytical methods used to derive our results proved to be highly precise, even when the NEW systematic uncertainty was treated as uncertain and the conditions for applying perturbative techniques, as outlined in Eq. (56), were not fully met. This is illustrated in Fig. 5, where we compare the analytical results obtained using the perturbative approach outlined in Sec. 4 at orders $\varepsilon_{s}^{2}$ with those from a numerical approach. This comparison focuses on the case where the NEW systematic uncertainty is treated as uncertain. The numerical results were obtained by numerically maximizing the likelihood over all parameters and calculating the expectation value of the profile likelihood ratio to determine the Bartlett correction, using MC simulations. Details of the simulation process are provided in App. C.
Figure 5: Comparison of analytical perturbative results at order $\varepsilon_{s}^{2}$ (orange line) and numerical results (blue line)
6 Conclusions
The Gamma Variance Model (GVM) provides a powerful statistical framework for addressing uncertainties in the assignment of systematic errors, informally referred to as errors-on-errors. Going beyond previous publications bib:Cowan2019 ; bib:Canonero2023 , we have extended the original model to include non-trivial correlations (not only 00 or ±$\pm 1$) and have provided formulas that greatly simplify its application in practice. As a by-product we have derived a useful connection between the BLUE method for combining results when there are non-trivial correlations and the corresponding likelihood method using nuisance parameters.
We applied the framework to the 7-8 TeV ATLAS-CMS top quark mass combination bib:ATLAS/CMS . All the results have been studied by considering various systematic uncertainties as uncertain, one at a time, and varying their associated error-on-error parameters $\varepsilon_{s}$. The aim of this paper has been to illustrate the impact of errors-on-errors on the combination, rather than assigning precise uncertainties to the systematic errors. This methodology can be used as a general procedure to identify the systematic uncertainties to which a combination is sensitive when they have associated errors-on-errors, and as a general approach to assess the robustness of a combination against systematic uncertainties that are themselves uncertain.
We conclude that the ATLAS-CMS $m_{t}$ combination is robust, with the central value remaining very stable across a broad spectrum of assumed values for errors-on-errors. The confidence interval remains stable as well, though it exhibits up to a $10\%$ increase when the b-JES systematic uncertainty is considered uncertain.
The scenario of an outlier present in the combination was also explored. This study is relevant to demonstrate the model’s properties in situations that may occur in future combinations. This example demonstrated the model’s sensitivity to the internal compatibility of the dataset. Specifically, the central value of the combination is less biased by outliers as the errors-on-errors parameters increase. Meanwhile, the confidence interval inflates with rising internal inconsistency among the data, as the GVM treats internal inconsistency in the input data as an additional source of uncertainty.
Acknowledgements
The authors are grateful to Mark Owen, Véronique Boisvert, and colleagues in the ATLAS Collaboration for their valuable feedback on this work. We also thank Richard Lockhart and Sara Algeri for organizing the Phystat review on the GVM, as well as Matthew Reader and Robert Thorne for productive discussions on the errors-on-errors framework. Finally, we acknowledge the support of the U.K. Science and Technology Facilities Council.
Data Availability Statement
This manuscript utilizes data from previously published studies, specifically in Sec. 5, where references are provided in the text. The remaining sections do not contain new data but are based on the results of calculations.
Appendix A Profiling over NPs modeling correlated auxiliary measurements
The BLUE log-likelihood, with the covariance matrix defined as in Eq.(8), can be derived by profiling Eq.(10) over all the NPs $\theta_{s}^{i}$ when the auxiliary measurements $u_{s}^{i}$ are set to 00. This process involves profiling the likelihood recursively over all $\theta_{s}^{i}$ for a given $s$. Here, we demonstrate how to do this for $s=M$. To simplify the notation we define
$\tilde{\mu}_{i}=\mu+\sum_{s=1}^{M-1}\Gamma_{i}^{s}\theta_{s}^{i}\,$ (39)
and
$K=-\frac{1}{2}\sum_{s=1}^{M-1}\sum_{i,j=1}^{N}\frac{\theta^{i}_{s}\theta^{j}_{s}}{\sigma_{u_{s}}^{2}}\left(\rho^{(s)}\right)_{ij}^{-1}\,.$ (40)
With these choices Eq.(8) becomes
$\begin{split}\ell(\mu,\boldsymbol{\theta})&=-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\tilde{\mu}_{i}-\Gamma_{i}^{M}\theta^{i}_{M}\right)V_{ij}^{-1}\\ &\times\left(y_{j}-\tilde{\mu}_{j}-\Gamma_{j}^{M}\theta^{j}_{M}\right)\\ &-\frac{1}{2}\sum_{i,j=1}^{N}\frac{\theta^{i}_{M}\theta^{j}_{M}}{\sigma_{u_{M}}^{2}}\left(\rho^{(M)}\right)_{ij}^{-1}+K\,.\end{split}$ (41)
To further simplify the derivation we define $\tilde{\theta}_{i}=\Gamma^{M}_{i}\theta^{i}_{M}$ and $\tilde{\rho}^{(M)}_{ij}=\rho^{(M)}_{ij}\Gamma^{M}_{i}\Gamma^{M}_{j}\sigma_{u_{M}}^{2}$. Thus the log-likelihood can be written in the simple form:
$\begin{split}\ell(\mu,\boldsymbol{\theta})=&-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\tilde{\mu}_{i}-\tilde{\theta}_{i}\right)V_{ij}^{-1}\left(y_{j}-\tilde{\mu}_{j}-\tilde{\theta}_{j}\right)\\ &-\frac{1}{2}\sum_{i,j=1}^{N}\tilde{\theta}_{i}\tilde{\theta}_{j}\left(\tilde{\rho}^{(M)}\right)_{ij}^{-1}+K\,.\end{split}$ (42)
To profile over the $\tilde{\theta}_{i}$, we need to solve the following equations for $\tilde{\theta}_{i}$ for all $i$:
$\frac{\partial\ell}{\partial\tilde{\theta}_{i}}=\sum_{j}^{N}(y_{j}-\tilde{\mu}_{j})V_{ij}^{-1}-\tilde{\theta}_{j}\left[V_{ij}^{-1}+\left(\tilde{\rho}^{(M)}\right)_{ij}^{-1}\right]=0\,.$ (43)
This leads to the profiled values of the $\tilde{\theta}_{i}$ parameters,
$\widehat{\widehat{\tilde{\theta}_{i}}}=\sum_{j,k}^{N}M^{-1}_{ij}V^{-1}_{jk}(y_{k}-\tilde{\mu}_{k})\,,$ (44)
where the matrix $M$ is defined by
$M_{ij}=V_{ij}^{-1}+\left(\tilde{\rho}^{(M)}\right)_{ij}^{-1}\,.$ (45)
This matrix is symmetric and we will use this property to switch its indices for the proof. If we now substitute the profiled values $\hat{\hat{\tilde{\theta}}}_{i}$ back into the log-likelihood we obtain
$\small\begin{split}&\ell(\mu,\boldsymbol{\theta_{1}},..,\boldsymbol{\theta_{M-1}},\boldsymbol{\widehat{\widehat{\tilde{\theta}}}_{M}})=\\ &-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\tilde{\mu}_{i}\right)\left[V_{ij}^{-1}-\sum_{p,q=1}^{N}V_{ip}^{-1}M^{-1}_{pq}V_{qj}^{-1}\right]\left(y_{j}-\tilde{\mu}_{j}\right)+K\,.\end{split}$ (46)
To further simplify this expression, the Woodbury identity bib:woodbury can be used to show that
$\begin{split}\left[V^{-1}-V^{-1}M^{-1}V^{-1}\right]_{ij}^{-1}&=V_{ij}-\left(-M+V^{-1}\right)_{ij}^{-1}\\ &=V_{ij}+\tilde{\rho}_{ij}^{(M)}\,.\end{split}$ (47)
Therefore we find
$\begin{split}\ell(\mu&,\boldsymbol{\theta_{1}},..,\boldsymbol{\theta_{M-1}},\boldsymbol{\widehat{\widehat{\tilde{\theta}}}_{M}})=\\ &-\frac{1}{2}\sum_{i,j=1}^{N}\left(y_{i}-\tilde{\mu}_{i}\right)\left[V+\tilde{\rho}^{(M)}\right]_{ij}^{-1}\left(y_{j}-\tilde{\mu}_{j}\right)+K\,,\end{split}$ (48)
where $\tilde{\rho}^{(M)}_{ij}=\rho^{(M)}_{ij}\Gamma^{M}_{i}\Gamma^{M}_{j}\sigma_{u_{M}}^{2}$ represents the $M$-th term in the systematic part of the BLUE covariance matrix $W$ as defined in Eq. (8), i.e., $U^{(M)}=\tilde{\rho}^{(M)}$. This procedure can be recursively applied to compute all terms in the BLUE covariance matrix corresponding to every source of systematic uncertainty. Thus, one obtains $W=V+\sum_{s=1}^{M}\tilde{\rho}^{(s)}=V+\sum_{s=1}^{M}U^{(s)}$.
Appendix B Useful formulas for the GVM with non-trivial correlations
In this appendix, we provide useful formulas to facilitate the implementation of the GVM with non-trivial correlations, similar to the approach outlined in Sec. 4 for the case of trivial correlations.
B.1 Profiled values of nuisance parameters
The method outlined in Sec. 4.1 for profiling the NPs can be extended to solve the score equations perturbatively for the GVM with arbitrary correlations, as defined in Eq. (16). Here we define the perturbative series for $\hat{\hat{\theta}}_{f(s,i)}$ as
$\hat{\hat{\theta}}_{f(s,i)}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\varepsilon_{s}^{4}\hat{\hat{\theta}}_{f(s,i)}^{(2)}+\cdots\,,$ (49)
where we introduce the composite index $f(s,i)=N(s-1)+i$ to denote $\theta_{s}^{i}$ for notational convenience. The first term in this series is given by
$\begin{split}\hat{\hat{\theta}}_{f(s,i)}^{(0)}&=\sum_{p=1}^{M}\sum_{j=1}^{N}\left(C^{(0)}\right)^{-1}_{f(s,i)\,f(p,j)}\left[\sum_{k=1}^{N}\Gamma_{j}^{p}V_{jk}^{-1}\left(y_{k}-\mu\right)\right.\\ &\left.+\sum_{k=1}^{N}\frac{\left(\rho^{(p)}\right)_{jk}^{-1}u_{p}^{k}}{v_{p}}\right]\,,\end{split}$ (50)
where the matrix $C^{(0)}_{f(s,i)\,f(p,j)}$ is defined as
$C^{(0)}_{f(s,i)\,f(p,j)}=\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{v_{s}}\left(\rho^{(s)}\right)_{ij}^{-1}\,.$ (51)
This first-order result is also the solution to the quadratic likelihood given by Eq. (10). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative term $\varepsilon_{s}^{2n}\,\hat{\hat{\theta}}_{f(s,i)}^{(n)}$ is given by
$\begin{split}&\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(n)}=\sum_{p=1}^{M}\sum_{j=1}^{N}\left(C^{(n)}\right)^{-1}_{f(s,i)\,f(p,j)}\\ &\times\left[\sum_{k=1}^{N}\Gamma_{j}^{p}V_{jk}^{-1}\left(y_{k}-\mu-\sum_{p^{\prime}=1}^{M}\Gamma_{k}^{p^{\prime}}\,T_{f(p^{\prime},k),n-1}\right)\right.\\ &\quad+\left.\frac{\sum_{k=1}^{N}\left(\rho^{(p)}\right)_{jk}^{-1}\left(u_{p}^{k}-T_{f(p,k),n-1}\right)}{S_{p,n-1}^{2}}\right]\,.\end{split}$ (52)
Here, the matrix $C^{(n)}_{f(s,i)\,f(p,j)}$ is defined as
$C^{(n)}_{f(s,i)\,f(p,j)}=\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{S_{s,n-1}^{2}}\left(\rho^{(s)}\right)_{ij}^{-1}\,,$ (53)
while $T_{f(s,k),n}$ and $S_{s,n}^{2}$ are defined as
$T_{f(s,k),n}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\cdots+\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(2n)}$ (54)
$\begin{split}S_{s,n}^{2}&=\frac{1}{1+2N\varepsilon_{s}^{2}}\left[v_{s}+2\varepsilon_{s}^{2}\sum_{i,j=1}^{N}\left(u_{s}^{i}-T_{f(s,i),n}\right)\right.\\ &\times\left(\rho^{(s)}\right)_{ij}^{-1}\left.\left(u_{s}^{j}-T_{f(s,j),n}\right)\right]\,.\end{split}$ (55)
Notice that if the condition
$\frac{2\varepsilon_{s}^{2}}{v_{s}}\sum_{i,j=1}^{N}\left(u_{s}^{i}-\hat{\hat{\theta}}_{f(s,i)}\right)\left(\rho^{(s)}_{ij}\right)^{-1}\left(u_{s}^{j}-\hat{\hat{\theta}}_{f(s,j)}\right)<1$ (56)
is not satisfied, the perturbative series is not guaranteed to converge.
B.2 Confidence intervals
We can compute the Bartlett correction for the GVM generalized to handle any correlations using the same approach described in Sec. 4.2. Specifically, for the likelihood defined by Eq. (16), the factor $b_{\mu}$ can be expressed as
$b_{\mu}=b_{\mu\boldsymbol{\theta}}-\tilde{b}_{\boldsymbol{\theta}}\,,$ (57)
where the terms $b_{\mu\boldsymbol{\theta}}$ and $\tilde{b}_{\boldsymbol{\theta}}$ are defined as
$\small\begin{split}b_{\mu\boldsymbol{\theta}}&=\sum_{s=1}^{M}\left\{4\frac{\text{Tr}\left[j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]}{\widehat{\sigma_{u_{s}}^{2}}}+\frac{\left(\text{Tr}\left[j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]\right)^{2}}{\left(\widehat{\sigma_{u_{s}}^{2}}\right)^{2}}\right.\\ &\left.-2\frac{\text{Tr}\left[j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]}{\left(\widehat{\sigma_{u_{s}}^{2}}\right)^{2}}\right\}\varepsilon_{s}^{2}\,,\end{split}$ (58)
$\small\begin{split}\tilde{b}_{\boldsymbol{\theta}}&=\sum_{s=1}^{M}\left\{4\frac{\text{Tr}\left[\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]}{\widehat{\sigma_{u_{s}}^{2}}}+\frac{\left(\text{Tr}\left[\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]\right)^{2}}{\left(\widehat{\sigma_{u_{s}}^{2}}\right)^{2}}\right.\\ &\left.-2\frac{\text{Tr}\left[\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]}{\left(\widehat{\sigma_{u_{s}}^{2}}\right)^{2}}\right\}\varepsilon_{s}^{2}\,.\end{split}$ (59)
Here, $\widehat{\sigma_{u_{s}}^{2}}$ is the MLE of $\sigma_{u_{s}}^{2}$, which can be computed by evaluating Eq. (15) at $\hat{\theta}_{s}^{i}$. The matrices $j^{-1}$ and $\tilde{j}^{-1}$ are defined by Eqs. (32) and (33), respectively, and are applied to the quadratic likelihood given in Eq. (10). Additionally, $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ and $\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ are submatrices of $j^{-1}$ and $\tilde{j}^{-1}$ that correspond to a fixed index $s$. Both $j^{-1}$ and $\tilde{j}^{-1}$ are functions of $\sigma_{u_{s}}^{2}$ and must be evaluated at $\widehat{\sigma_{u_{s}}^{2}}$. To simplify the computation of the Bartlett correction, the double-indexed NPs $\theta_{s}^{i}$ can be flattened into single-indexed variables by defining $f(s,i)=N(s-1)+i$, so that $\theta_{s}^{i}=\theta_{f(s,i)}$.
B.3 Goodness-of-fit
The GOF statistic of the GVM with non-trivial correlations (Eq. (16)) is defined as
$\begin{split}q&=\sum_{i,j=1}^{N}\left(y_{i}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{i}^{s}\hat{\theta}^{i}_{s}\right)V_{ij}^{-1}\left(y_{j}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{j}^{s}\hat{\theta}^{j}_{s}\right)\\ &+\sum_{s=1}^{M}\left(N+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+\frac{2\varepsilon_{s}^{2}}{v_{s}}\sum_{i,j=1}^{N}(u^{i}_{s}-\hat{\theta}^{i}_{s})\right.\\ &\times\left.\left(\rho^{(s)}\right)_{ij}^{-1}(u^{j}_{s}-\hat{\theta}^{j}_{s})\right]\,.\end{split}$ (60)
This GOF statistic can be corrected using the Bartlett correction as outlined in Sec. 4.3. Specifically the Bartlett factor is equal to
$b_{q}=(2N+N^{2})\sum_{s=1}^{M}\varepsilon_{s}^{2}-b_{\mu\boldsymbol{\theta}}\,,$ (61)
where the definition of $b_{\mu\boldsymbol{\theta}}$ is given in Eq. (58).
Appendix C Generating toy data
To estimate the expectation values of the profile likelihood ratio and the GOF with enhanced precision toy data can be generated by fixing the parameters at their MLE values, $\hat{\mu}$ and $\boldsymbol{\hat{\theta}}$. When using the GVM likelihoods, it is also necessary to fix the parameters $\sigma_{u_{s}}^{2}$ to some value. While one could use the MLE of $\sigma_{u_{s}}^{2}$, this estimator is biased, and for large $\varepsilon_{s}^{2}$, it may lead to imprecise results. Instead, it is preferable to use an unbiased estimator at order $\varepsilon_{s}^{2}$. For Eq. (11), the unbiased estimator is defined as
$\widehat{\sigma_{u_{s}}^{2}}^{\ast}=\frac{v_{s}+2\varepsilon_{s}^{2}(u_{s}-\hat{\theta}_{s})^{2}}{1+2\varepsilon_{s}^{2}}+2\varepsilon_{s}^{2}j^{\theta_{s}\theta_{s}}\,,$ (62)
where $j^{\theta_{s}\theta_{s}}$ is computed as in Eq. (31). For Eq. (14), the unbiased estimator becomes:
$\begin{split}\widehat{\sigma_{u_{s}}^{2}}^{\ast}&=\frac{v_{s}+2\varepsilon_{s}^{2}\sum_{i,j}^{N}(u_{s}^{i}-\hat{\theta}_{s}^{i})\left(\rho^{(s)}\right)_{ij}^{-1}(u_{s}^{j}-\hat{\theta}_{s}^{j})}{1+2N\varepsilon_{s}^{2}}\\ &+2\varepsilon_{s}^{2}\text{Tr}\left[j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}\left(\rho^{(s)}\right)^{-1}\right]\,,\end{split}$ (63)
where $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ is computed as in Eq. (58). These unbiased estimators were computed following the approach described in bib:Cox1968 ; bib:Codeiro1991 ; bib:Codeiro2014 . They are then used to generate auxiliary measurements, as well as the stochastic variables $v_{s}$, which follow a Gamma distribution:
$v_{s}\sim\frac{\beta_{s}^{\alpha_{s}}}{\Gamma(\alpha_{s})}v_{s}^{\alpha_{s}-1}e^{-\beta_{s}v_{s}}\,,$ (64)
where $\alpha_{s}=\frac{1}{4\varepsilon_{s}^{2}}$ and $\beta_{s}=\frac{1}{4\sigma^{2}_{u_{s}}\varepsilon_{s}^{2}}$.